### v1

# CONLINE: Complex Code Generation and Refinement with Online Searching and Correctness Testing

Abstract
Large Language Models (LLMs) have revolutionized code generation ability by converting natural language descriptions into executable code. However, generating complex code within real-world scenarios remains challenging due to intricate structures, subtle bugs, understanding of advanced data types, and lack of supplementary contents. To address these challenges, we introduce the CONLINE framework, which enhances code generation by incorporating planned online searches for information retrieval and automated correctness testing for iterative refinement. CONLINE also serializes the complex inputs and outputs to improve comprehension and generate test case to ensure the framework’s adaptability for real-world applications. CONLINE is validated through rigorous experiments on the DS-1000 and ClassEval datasets. It shows that CONLINE substantially improves the quality of complex code generation, highlighting its potential to enhance the practicality and reliability of LLMs in generating intricate code.
CONLINE: Complex Code Generation and Refinement
with Online Searching and Correctness Testing
Xinyi He1thanks: The contributions by Xinyi He, Jiaru Zou and Yun Lin have been conducted and completed during their internships at Microsoft.   Jiaru Zou2*   Yun Lin3*   Mengyu Zhou4thanks: Corresponding author.   Shi Han 4   Zejian Yuan1   Dongmei Zhang4 1 Xi’an Jiaotong University 2 University of Illinois at Urbana-Champaign 3 Peking University 4 Microsoft Research hxyhxy@stu.xjtu.edu.cn, linyun@stu.pku.edu.cn, jiaruz2@illinois.edu, yuan.ze.jian@xjtu.edu.cn, {mezho, shihan, dongmeiz}@microsoft.com
1 Introduction
Figure 1: An Example of the Human Programmer Code-writing Process Imitated by the CONLINE. After the problem is received, an online search is performed to simulate search results and create an initial version of code. Test cases are then generated, and the code is executed to produce output results. The code is refined based on the correctness of these results.
Automated code generation from natural language descriptions has became more realistic, as large language models (LLMs) show their potential to generate complex and accurate code Manna and Waldinger (1971); Li et al. (2023); Luo et al. (2023); Rozière et al. (2024). There are a variety of methods that have been proposed to improve the quality of LLM code generation, such as retrieving offline documents Zhou et al. (2023); Jiang et al. (2023) and debugging generated code Zhang et al. (2023); Chen et al. (2023).
In real-world scenarios, complex code generation can involve intricate problem description, sophisticated code logic, and advanced data types Lai et al. (2022); Du et al. (2023); He et al. (2023). For example (see Figure 1), solving the required code generation involves multiple transformations of various matrices, as well as the invocation of functions that demands numerous and stringent parameters. Existing methods were not enough to address the following new challenges that emerge during complex code generation:
  1. 1.
    Complex code generation is hindered by intricate structures and logic. E.g., the complexity of the example in Figure 1 is as previously described.
  2. 2.
    Obscure bugs in complex code necessitate critical detection and refinement. Existing methods often refine code based on execution errors testing Zhang et al. (2023); Jiang et al. (2023), but for complex codes, sometimes the executable code has non-obvious bugs,e.g., initial code in Figure 1.
  3. 3.
    The complexity of inputs and outputs hampers LLM’s comprehension ability during code refinement. In complex code, inputs and outputs like Pandas DataFrames, PyTorch tensors, and matplotlib PNG images are intricate. Understanding such inputs and outputs poses challenges for LLMs due to large matrices, image modalities and so on.
  4. 4.
    Real-world scenarios frequently lack supplemental content, such as test cases and documentation pools. Existing works predominantly rely on pre-existing test cases in datasets Zhang et al. (2023); Jiang et al. (2023) or off-line document pools Zhou et al. (2023); Jiang et al. (2023), which are challenging to obtain directly in real-world scenarios.
To address the above challenges, we propose a new code generation framework CONLINE that enhances the code generation and refinement quality of LLMs with planned online searching and automated correctness testing steps in this paper. The intuition of CONLINE is straightforward: During the process of coding, most human programmers are not bothered by the above challenges as shown in Figure 1. They could easily bypass them by searching on-line for answers, experiences, or guidelines. Also, they are able to design test cases and run code against them to verify the correctness of the code logic.
Correspondingly, CONLINE uses online search for code generation to partially address Challenges 1 and 4. For Challenge 1, it offers following advantages: (1) Retrieving information from library documents allows the retrieval of standardized function syntax, preventing errors from model hallucinations e.g., generating non-existing parameters. (2) Retrieving information from the latest blogs or Q&A platforms, such as StackOverflow, facilitates the emulation of commonly used code patterns, thereby reducing the complexity of generated code. For Challenge 4, online search encompasses the content of offline document pools and extends beyond with its superior capability to adapt to the diverse and evolving needs of real-world scenarios. Directly searching for solutions to such problems is difficult, because the problems are intricate and may involve multiple challenges. Therefore, we propose online search with generation query through planning.
To address the Challenge 2, this work proposes correctness testing in refinement. LLMs may generate complex code that contains subtle errors, resulting in code that is executable but does not produce the desired output. Therefore, during the refinement process, it is more critical to verify that the executed code produces the correct results rather than just checking for the existence of the errors. We propose incorporating both the execution output results and the errors as integral parts of the refinement to enhance the correctness of the code.
In refinement with correctness testing, we propose generation of test cases to address the Challenge 3. Test cases are crucial, as they serve as indispensable inputs for the code execution. This work utilizes LLMs to automatically generate test cases, making the framework applicable even in real-world scenarios where no known test cases exist. To address the Challenge 4, we propose serialization of input and output. It renders the input and output more intuitive and understandable for the model, particularly for those that are excessively long or non-textual modalities.
We test the effectiveness of CONLINE on two complex code generation datasets using three different LLMs. For the complex code generation datasets, we selected DS-1000, which focuses on complex data science code generation, and the Class Eval dataset, which is oriented towards class-level code generation. By comparing with the existing state-of-the-art (SOTA) baseline, we achieved a 10.9% improvement on DS-1000 and average 9.47% on ClassEval, confirming the effectiveness of CONLINE. Through comprehensive ablation studies, we verified the effectiveness of each component of CONLINE. Moreover, given the varying performance of CONLINE on different base models, we discovered that CONLINE requires models to have capabilities such as Planning, which vary according to the complexity of the problem, in order to further enhance the model’s code generation abilities.
2 Related Work
The realm of automated code generation has been propelled by benchmark datasets such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and APPS Hendrycks et al. (2021), which assess the proficiency of language models in generating executable code from descriptions. These datasets encompass a variety of programming problems, yet recent studies have sought to escalate the complexity of code generation tasks. Works like DS-1000 Lai et al. (2022), ClassEval Du et al. (2023) and Text2Analysis He et al. (2023) have introduced datasets targeting specialized domains, including data science, object-oriented class generation, and data analysis. These endeavors reflect an emerging trend towards enhancing models’ abilities to produce sophisticated and domain-specific code structures.
Recent research has concentrated on improving Large Language Models (LLMs) for code generation via retrieve, refine and so on strategies. Retrieve enhances model performance by pulling in external information, with techniques like DocPrompt Zhou et al. (2023) accessing fixed document repositories and SELVEVOLVE Jiang et al. (2023) using models as knowledge bases. However, these are limited to document-sourced content and provide limited benefits.
Refine iteratively enhances generated code for greater precision. Projects such as Self-Debug Chen et al. (2023) and SELFEVOLVE Jiang et al. (2023) focus on error refinement, while Self-EditZhang et al. (2023) uses code execution results as feedback. Despite advancements, these methods often miss subtle bugs and are less applicable to complex, real-world code generation without readily available test cases or ground truths.
Figure 2: The Pipline of CONLINE. Step 1: LLM is employed to strategize the Problem and formulate queries based on the outlined steps. These queries enable the retrieval of diverse information from the internet. A high-quality initial code can be obtained through effective planning and leveraging internet information. Step 2: LLM generates test cases for testing the initial code. The test results serve as crucial inputs for the subsequent cycle of code refinement. Through iterative refinement processes, the quality of the initial code can be significantly improved.
3 Methodology
The code generation task involves predicting a solution code $W$ given a problem description $D$. When given an input $i$, the execution of code $W$ produces an output result $o$ and a potential error $e$, where both $o$ and $e$ can be empty $\varnothing$. The generated codes are evaluated against a set of test cases and ground truth $\{(t_{j},g_{j})\}_{j=1}^{J}$. The correctness of the code $W$ is determined by verifying $o_{j}=g_{j}\land e_{j}=\varnothing$ when all $i_{j}=t_{j},j\in\{1,\ldots,J\}$.
In this work, we adopt a two-step approach for code generation, mirroring the way humans write code. The first step is retrieve, where relevant information is obtained through online saerch and utilized by LLMs to generate initial code. The second step is refine, where the initial code is refined based on the execution results, leading to the generation of the final version of the code.
3.1 Retrieve
The difficulty in achieving effective online retrieval lies in formulating optimal search queries. On the one hand, for complex code generation, the problems are intricate and may involve multiple challenges. Directly searching for solutions to such problems is inaccurate and difficult. On the other hand, it is challenging that matching queries directly through methods for offline document like similarity calculations, due to the nature of online libraries. Therefore, we propose using planning to solve the challenge.
The retrieval process is divided into three steps: 1. Search queries $Q=\{q_{1},\ldots,q_{N}\}$ are generated through planning. 2. Conducting online searches using these queries to obtain relevant background information $INFO=\{info_{1},\ldots,info_{M}\}$. 3. The initial code $W_{0}$ is generated by the LLMs $\theta$ with the information obtained $I$:
$\widehat{W}_{0}\sim p_{\theta}(.|D,I)$ (1)
3.1.1 Generation Query through Planning
In order to generate more targeted queries, we initiate the process by using LLMs to do planning regarding the given problem. The planning phase involves outlining the natural language steps $P=\{plan_{1},\ldots,plan_{N}\}$ required to address the problem. Later, the assessment involves utilizing LLMs to determine whether each planning step requires an online search. Subsequently, the planning steps identified as necessitating online search are translated into queries $Q=\{q_{1},\ldots,q_{N}\}$ for use in the subsequent search process.
$\widehat{P},\widehat{Q}\ \sim p_{\theta}(.|D)$ (2)
3.1.2 Online Search
For the above generated queries, we conduct an online search. We use the online search API [^1] https://github.com/Nv7-GitHub/googlesearch for the search process.
$\{url_{1},\ldots,url_{N_{u}}\}=search(q_{j}),j\in\{1,\ldots,N_{q}\}$
where, $N_{q}$ is the number of queries for the problem, $N_{u}$ is the number of urls for one query. In this study, we use $N_{q}=1$ and $N_{u}=1$.
Through the analysis of the website distribution Table 4, we observed that more than 90% of the URLs are concentrated on a total of 8 websites. Specific extraction rules are established for prominent websites such as StackOverflow to extract key information, facilitating a more comprehensive understanding of the website’s content by subsequent models. Generic extraction rules are employed for extracting key information from other websites.
$info_{j,k}=extract(url_{k}),k\in\{1,\ldots,N_{u}\}$
The information $INFO$ is composed of details from each query $q_{j}$, each url $url_{k}$, with each piece of information $info_{j,k}$ extracted.
3.2 Refine
Existing work Chen et al. (2023); Jiang et al. (2023) typically emphasizes correction of errors identified during the refinement process. However, we observe that refining code that produces error-free outputs is equally crucial during the refinement process. Therefore, we introduce correction testing in §3.2.1. Additionally, we propose methods for the generation of test cases and serialization of inputs and outputs during the refinement process.
3.2.1 Correctness Testing
Correctness testing refers to the refinement of generated code based on correctness, determined by analyzing errors and output results obtained during code execution. In the context of complex code generation, the intricate logic of the code makes it challenging for the LLMs to consider every detail during code generation, and precisely ascertain the results obtained at each step of the execution process. Consequently, some code may be executed without errors, producing output results that do not align with what is expected. Incorporating both the error and the output result into the refinement process allows the model to take advantage of self-correction mechanisms.
$\left\{\begin{array}[]{ccl}e_{j,k},o_{j,k}&=&execute(W_{j},i_{k}),j\in\{1,\ldots,N_{f}\}\\ INFO_{e_{j,k}}&=&\{e_{j,k},extract(search(e_{j,k}))\}\\ \widehat{W}_{j+1}&\sim&p_{\theta}(.|D,W_{j},\{S_{i},S_{o_{j}},INFO_{e_{j}}\}_{k}),\\ &&\hfil k\in\{1,\ldots,N_{i}\}\end{array}\right.$
where, $N_{f}$ is the total number of refinement steps, $N_{i}$ is the number of inputs. $i_{K}$ is the k-th input for the problem from Equation (3), $S_{i}$ and $S_{o_{j}}$ is the serialization of the input and output from Equation (4).
3.2.2 Generation of Test Case
Test cases is crucial, as they serve as indispensable inputs for the code execution in refinement. While, existing works in refining code predominantly rely on pre-existing test cases in datasets Zhang et al. (2023); Jiang et al. (2023), which are challenging to obtain directly in real-world scenarios. Moreover, some existing work Chen et al. (2023) even uses the ground truth output of the test case to refine the code, which is even more challenging to obtain for complex code problems in real-world scenarios. Because their problems involve various logical operations, and even with input test cases, deriving answers directly without code-based computations is demanding.
In this study, we propose generation of test case with LLMs to adapt to real-world scenarios.
$\left\{\begin{array}[]{ccl}\widehat{I}&\sim&p_{\theta}(.|D)\\ I&=&\{i_{1},\ldots,i_{N_{i}}\}\end{array}\right.$ (3)
3.2.3 Serialization of Input and Output
Serialization of input and output makes them more intuitive and understandable for the model. For complex code, some inputs and outputs are intricate, such as Pandas DataFrames, PyTorch tensors, and matplotlib PNG images. Understanding such inputs and outputs poses challenges for LLMs due to large matrices, image modalities, and so on.
In this study, we serialize common data structures in Python as follows:
  1. 1.
    For NumPy arrays, Pandas DataFrames, PyTorch tensors, and TensorFlow tensors, the serialization includes a truncated string of data, data type, data shape, and statistical information.
  2. 2.
    For image structures (such as PNG images generated by the Matplotlib library), we serialize them into SVG (Scalable Vector Graphics) format for LLMs to comprehend.
$S_{n}=serialize(n),n\in\{i_{k},o_{j,k}\}$ (4)
Table 1: Table of Main Results and Ablation Study for DS-1000. The base model for CONLINE is GPT-4. All metric numbers are represented as percentages. For each metric, the bold number indicates the highest performance.
Method Perturbation Total/Avg.
Origin Surface Semantic Diff-Rewrite
DocPrompting 53.95 50.00 38.39 21.05 43.30
Self-Debugging 63.38 59.21 45.65 28.40 53.00
SELFEVOLVE 66.23 67.11 48.70 33.95 57.10
CONLINE 71.71 74.34 66.96 53.09 68.00
   w/o refinement of output 68.42 69.74 62.61 48.77 64.10
   w/o refinement of error 68.20 73.03 62.61 49.38 64.60
   w/o serialization 70.18 75.00 65.22 51.23 66.70
   w/o generation of test case 66.23 71.05 59.57 45.68 62.10
   w/o online retrieval 68.64 70.39 60.00 51.23 64.10
   GPT-4 64.47 69.74 56.96 43.83 60.20
4 Experiment
4.1 Experiment Setup
4.1.1 Datasets
We conduct experience on two complex code generation datasets:
1. DS-1000 Lai et al. (2022): DS-1000 is a code generation benchmark with a thousand data science questions spanning seven Python libraries. The complexity of this dataset is manifested in two aspects. First, complexity arises from intricate logical reasoning required during code generation due to the complex nature of the problems. For example, on the DS-1000 dataset, the average length of problem descriptions is 140 words, whereas other commonly used code generation datasets such as HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) have lengths of 23 and 15.7 words, respectively. Secondly, the input-output involves various complex data structures related to data science, making the code logic intricate during transformations of the data.
The dataset provides both Insertion and Completion style prompts, where the data is the same, differing only in prompt format, thus yielding similar results. In this paper, experiments are conducted with the Completion style prompt.
2. ClassEval Du et al. (2023): ClassEval is the first class-level Python code generation benchmark designed to evaluate code generation models’ performance on a diverse set of object-oriented programming tasks. The dataset comprises a curated collection of 100 tasks. These tasks cover a wide range of concepts, including inheritance, polymorphism, encapsulation, etc. Each coding task is in the format of the class skeleton, outlining target method description inside the class. The complexity of this dataset resides in its abstraction and hierarchical class structure. Tested models must generate large-scale code units and establish connections between each target method within the entire class, rather than focusing solely on individual functions.
The dataset provides two prompt designs for LLMs with or without IF ability. In our experiments, we employ the class skeleton as the prompt for GPT-based models, a system prompt along with task instructions for the WizardCoder.
4.1.2 Evaluation
We employ the same evaluation methodology as the original datasets for both DS-1000 and ClassEval.
  • DS-1000. we follow the original dataset using Pass@1 accuracy. This evaluation was conducted across total and four types of perturbations: Origin, Surface, Semantic, and Diff-Rewrite.
  • ClassEval. we follow the original dataset and evaluate model performance through Pass@$K$ metric. Specifically, we calculate both class-leval and method-level Pass@$K$ with $K=1,3,5$.
4.1.3 Base LLMs
This work primarily utilizes the GPT OpenAI (2023) series as the LLM base model to validate the effectiveness of the framework. GPT-4 is utilized in its gpt-4-32k-0613 version, while GPT-3.5 is utilized in the gpt-35-turbo-16k-0613 version. To further investigate the performance of CONLINE on both open-source and specialized code generation models, we have also employed WizardCoder Luo et al. (2023) as a base model with WizardCoder-Python-13B-V1.0 version.
4.1.4 Baselines
For the DS-1000, we selected three LLM-based frameworks as baselines.
1. DocPrompt Zhou et al. (2023): DocPrompting enhances the LLM by employing a fine-tuned retriever to fetch problem-relevant documentation from offline document pools. The model then conditions on these documents, along with the problem description, to generate code.
2. Self-Debugging Chen et al. (2023): This approach depends on a SQL application or Python interpreter to instruct language models in revising SQL commands or Python code containing errors. For the sake of a fair comparison, we utilize its "simple" variant.
3. SELFEVOLVE Jiang et al. (2023): Employs LLMs as both sources of knowledge and self-reflective programmers. During the self-reflective process, it refines the code by addressing bugs.
For the ClassEval, We select four LLM-based code generation models as baselines: Instruct-CodeGen [^2] https://huggingface.co/sahil2801/instruct-codegen-16B, SantaCoder Allal et al. (2023), Instruct-StarCoder [^3] https://huggingface.co/GeorgiaTechResearchInstitute/ starcoder-gpteacher-code-instruct , WizardCoder Luo et al. (2023)
It is important to note that the test cases involving the refinement process in the baselines mentioned above all use the test cases from the dataset designated for testing. However, within the context of the real-world scenario of CONLINE, test cases from the dataset should not be used within the framework. Without these test cases, they are entirely incapable of functioning, hence we have positioned our CONLINE at a disadvantage for comparison purposes. Despite this, we are still able to achieve SOTA performance.
4.2 Main Results
Regarding the DS-1000 dataset, the main results are shown in Table 1. CONLINE surpasses the current SOTA model, SELFEVOLVE, by 10.9%, establishing itself as the new SOTA. Especially under the Diff-Rewrite perturbation setting, CONLINE exceeds SELFEVOLVE by 19.95%, which demonstrates the effectiveness of CONLINE in generating complex code. CONLINE employs online search and correctness testing to allow the model to imitate existing code patterns, thereby reducing the difficulty of generating new code, and refines the details to further enhance the correctness of the code.
Table 2: Table of Main Results and Ablation Study for ClassEval. All metric numbers are represented as percentages. For each metric, the bold number indicates the highest performance.
Method Class-level Method-level
Pass@1 Pass@3 Pass@5 Pass@1 Pass@3 Pass@5
Instruct-StarCoder 10.2 12.7 14.0 23,1 26.5 27.7
SantaCoder 8.6 9.9 10.0 27.7 33.0 34.9
Instruct-CodGen 8.2 12.3 13.0 24.9 34.3 37.1
WizardCoder 12.2 20.0 23.0 35.2 47.1 51.1
CONLINE (GPT-4) 46.3 49.5 52.8 67.9 72.5 77.6
   w/o refinement of output 43.5 46.8 51.4 66.4 69.0 73.4
   w/o refinement of error 46.2 49.5 51.7 67.9 72.5 77.2
   w/o generation of test case 42.7 47.9 50.6 65.9 70.8 72.4
   w/o online retrieval 37.2 42.5 44.9 60.4 65.7 69.8
   GPT-4 36.2 39.3 43.5 58.6 64.9 67.3
CONLINE (GPT-3.5) 33.9 43.6 45.8 58.4 62.0 64.7
   w/o refinement of output 31.2 40.7 43.5 56.8 60.3 64.2
   w/o refinement of error 30.7 42.5 41.4 53.6 62.0 61.3
   w/o generation of test case 31.0 40.2 41.9 55.4 59.1 61.7
   w/o online retrieval 31.4 40.6 42.8 55.7 59.4 62.3
   GPT-3.5 27.0 32.1 35.4 52.3 54.6 59.4
For the ClassEval dataset, the results are shown in Table 2. Our experiments demonstrate that CONLINE has an overall higher porformance on both class-level and method-level Pass@$K$ evaluation. Specifically, COLINE(GPT-4) outperforms the WizardCoder (best baseline model) significantly by an average of 31.3% on Class-level and 28.3% in Method level; COLINE(GPT-3.5) outporforms the WizardCoder by an average of 22.7& in Class-level and 17.2% in Method-level.
4.3 Ablation Study
In this work, to validate the effectiveness of CONLINE, we conduct different ablation studies, with results presented in Table 1 and Table 2. Details on the ablation study are shown in §C.
CONLINE significantly enhances the base model’s ability to generate complex code. Compared to the base model, CONLINE (GPT-4) has shown improvements of 7.8% on the DS-1000 dataset and average 9.47% on ClassEval, demonstrating the effectiveness of the CONLINE.
Online search, generation of test cases, and serialization each contribute to the model’s performance improvements. Compared to CONLINE, after performing ablation studies, these features showed a decrease in performance of 3.9%, 5.9%, and 1.3% respectively on the DS-1000 dataset. The online search improves the model by providing common code patterns, which reduces the difficulty of the model in generating initial code. Serialization, by converting inputs and outputs into a sequential format, allows the model to more intuitively observe inputs and outputs that are too lengthy or are in non-textual modalities, thereby strengthening its ability to solve complex code problems.
During the refinement process, correctness testing is crucial, meaning that both the output result and error are equally important. After separately conducting ablation studies on the output result and error, CONLINE shows a decrease of 3.9% and 3.4% respectively on the DS-1000 dataset, and average 2.7% and 0.3% on the ClassEval datset. This indicates that the output result contributes more to the refinement process than the error. However, in previous works, the output result is often overlooked, which should not be the case, especially in the generation of complex code. The evidence from the ablation study emphasizes the necessity of paying attention to the output results during the refinement phase to ensure the generation of high-quality, complex code.
4.4 Analysis of Different Base Models Performance
Table 3: Table of Different Base Models Results for DS-1000. All metric numbers are represented as percentages. For each metric in each section, the bold number indicates the highest performance.
Method Perturbation Total/Avg.
Origin Surface Semantic Diff-Rewrite
GPT-4 64.47 69.74 56.96 43.83 60.20
   w/ retrieve 66.23 71.05 59.57 45.68 62.10
   w/ refine 68.64 70.39 60.00 51.23 64.10
   CONLINE 71.71 74.34 66.96 53.09 68.00
GPT-3.5 57.02 43.42 40.00 32.72 47.10
   w/ retrieve 47.15 25.00 36.96 25.31 37.90
   w/ refine 55.70 50.66 44.35 35.80 49.10
WizardCoder 41.01 21.71 31.74 16.05 31.90
   w/ retrieve 15.79 9.21 12.17 9.88 13.00
   w/ refine 39.69 21.71 30.00 15.43 30.80
Table 3 shows the performance results of CONLINE on the DS-1000 dataset with different base models. We can see that GPT-4 has been comprehensively improved with CONLINE, but the performance on GPT-3.5 and WizardCoder is mixed. This indicates that CONLINE requires the model to have the following capabilities to enhance its performance:
For code generation planning ability, the higher the complexity of the code that needs to be generated, the higher the demand for planning ability. Planning capability is key to online retrieval; only correct planning can generate appropriate queries to retrieve useful information. After incorporating online retrieval, GPT-3.5 has an increase of 4.75% on ClassEval as shown in Table 2, yet it decreased by 9.2% on DS-1000 as shown in Table 3. For ClassEval, we plan for each function during online retrieval. The challenge of ClassEval lies in how to generate the entire class and the interrelated functions, but the complexity of individual function codes is not as high as that of DS-1000. Therefore, the planning ability of GPT-3.5 can handle ClassEval, but it is inferior on DS-1000.
Code generation also necessitates the models to possess the capability of in-context learning. The generated code should be built on all the above provided contents, and the understanding of the generated results based on the preceding input prompt is of great importance in the refinement stage. In Table 3, it is observed that WizardCoder has a noticeable drop of 18.9% and 1.1% on the DS-1000 dataset when utilizing online retrieval and refinement respectively. And the overall performance of WizardCoder on DS-1000 is comparatively interior to GPT models. One possible reason for this lies in the limited in-context learning capability of WizardCoder given some of prompts of DS-1000 are lengthy and complicated, which makes it difficult for WizardCoder to comprehend the related context and make corresponding code modifications accurately.
4.5 Case Study
Figure 3: Case Study for Correctness Testing.
For the case study on online retrieval, refer to Figure 4. It can be observed that by imitating the usage of functions found through online search, the model is better equipped to prepare the required parameters for the functions and to generate corresponding code. This significantly reduces the difficulty of generating complex code.
For the case study on correctness testing, refer to Figure 3. It is evident that, although the initially generated code did not show obvious errors, the output of the code did not align with the expected results. The model refines the code based on the output, thereby improving subtle errors and generating the correct code.
5 Conclusion
In conclusion, CONLINE introduces a novel framework for generating complex code in real-word scenarios by emulating human coding processes like online searching and test case creation. It effectively overcomes challenges in code structure and logic, subtle bug detection, and handling of complex data. The framework’s innovative use of online search, planning for query generation, correctness testing, and input-output serialization significantly improves code accuracy and model understanding. Tested on various datasets, CONLINE outperforms existing methods, demonstrating its efficacy in real-world code generation tasks.
Limitations
The main limitation of our research is that it has underlying issues of exceeding the allowed times of accesses due to multiple calls to the Google Search API. Similarly, we also have made multiple API calls to test and enhance the performance of the GPT models.
Ethics Policy
This research does not pose any ethical concerns. The datasets and other associated resources utilized in this study are publicly available and widely used in various other existing work.
Appendix A Details for on-line search
Table 4: Website Base Station Distributions Table During the DS-1000 Online Retrieval Process.
Website Proportion
https://stackoverflow.com 57.92%
https://numpy.org 8.59%
https://pandas.pydata.org 5.70%
https://www.geeksforgeeks.org 5.07%
https://docs.scipy.org 4.76%
https://matplotlib.org 3.20%
https://www.tensorflow.org 3.04%
http://scikit-learn.org 2.42%
The Others 9.29%
Appendix B Details for Base Models
The parameter details for each model in the experiment are as follows:
  • GPT-4: model: gpt-4-32k-0613, temperature: 0, top_p: 0.95, max_tokens: 1024.
  • GPT-3.5: model: gpt-35-turbo-16k-0613, temperature: 0, top_p: 0.95, max_tokens: 1024.
  • WizardCoder: WizardCoder-Python-13B-V1.0, temperature: 0, top_p: 0.95, max_tokens: 1024.
Appendix C Details for Ablation Study
  • Without refinement of output: During the refinement process, the output result is not refined; that is, refinement is conducted solely based on the error.
  • Without refinement of error: During the refinement process, the error is not refined; that is, refinement is conducted solely based on the output result.
  • Without serialization: During the refinement process, the input and output are not serialized; instead, their printout results are directly used as input.
  • Without generation of test cases: Test cases are not generated. Since refinement cannot be performed without test cases, only online retrieval is conducted.
  • Without online retrieval: Online retrieval is not performed, and the process is limited to refinement with correctness testing.
Appendix D Case Study
Figure 4 demonstrates an example of online retrieval process in DS-1000 dataset. Given a problem description, the tested model firstly attempts to use the tensor_scatter_nd_update function from TensorFlow to achieve the solution, but encounters difficulty in selecting the appropriate class indices. After unsuccessful attempts, the model turns to an online search through a self-generated query and integrating the research results as part of next prompt. Through the newfound knowledge online, the model re-implements a solution that efficiently generates the desired tensor representing the accuracy of each class. Overall, the process highlights how online retrieval can provide useful insights and potential solutions to complex challenges for LLMs’ code generation.
Figure 4: Case Study for Online Retrieval.
Appendix E Prompts of CONLINE
Figure 5: Plan and Queries Generation Prompt on DS-1000.
Figure 6: Plan and Queries Generation Prompt on ClassEval.
Figure 7: Online Retrieval Code Generation Prompt on DS-1000 and ClassEval.
Figure 8: Generation of Test Case Prompt on DS-1000.
Figure 9: Generation of Test Case Prompt on ClassEval.
Figure 10: Refinement with Correctness Testing Code Generation Prompt on DS-1000 and ClassEval.